several students form a circle by standing next to each other. samantha and aaron are directly across form each other and are approximately 30 feet apart. if each student takes up 4 feet of the circle, how many students are in the circle?
Answer by checkley77(12844)

CIRCUMFERENCE=2*PI*R
C=2*3.14*30/2
C=6.28*15
C=94.2
94.2/4=23.55 OR 24 STUDENTS.